OpenStudy (bringmethehorizon_):

WILL GIVE MEDAL!!!! Given the following functions f(x) and g(x), solve f[g(6)] and select the correct answer below: f(x) = 6x + 12 g(x) = x − 8 A: -96 B: 0 C: 24 D: 48

Vocaloid (vocaloid):

if g(x) = x − 8, what is g(6)?

OpenStudy (bringmethehorizon_):

10?

Vocaloid (vocaloid):

i'm not sure how you got that

OpenStudy (kgrendel0324):

14?

Vocaloid (vocaloid):

think carefully

OpenStudy (bringmethehorizon_):

8x6= 48 ?

Vocaloid (vocaloid):

g(x) = x - 8 means take the number inside the parentheses and subtract 8 so what is g(6)?

OpenStudy (kgrendel0324):

-2

Vocaloid (vocaloid):

@bringmethehorizon_ with me so far?

OpenStudy (bringmethehorizon_):

Yess I understand now, okay

Vocaloid (vocaloid):

good so g(6) = -2 now we find f(-2)

Vocaloid (vocaloid):

if f(x) = 6x + 12 what is f(-2) = ?

OpenStudy (bringmethehorizon_):

4?

OpenStudy (bringmethehorizon_):

wait

Vocaloid (vocaloid):

f(x) = 6x + 12 = 6(-2) + 12 = ?

OpenStudy (bringmethehorizon_):

I did 16 instead lol

OpenStudy (bringmethehorizon_):

0?

Vocaloid (vocaloid):

good, so 0 is your answer

OpenStudy (bringmethehorizon_):

Thank you :)